OpenStudy (anonymous):

MEDAL AND FAN: NEED SOME HELP!!! PLEASE!!! What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (tkhunny):

Have you considered subtracting one from the other?

OpenStudy (anonymous):

I'm confused on how to do this, please help?

OpenStudy (welshfella):

subtract term by term x - x = 0 so we ewliminate the x

OpenStudy (welshfella):

x - y = -3 x + 3y = 5 x-x = 0 -y - 3y = ? -3 - 5 = ?

OpenStudy (welshfella):

that will leave with an equation in one unknown (y) which you can easily solve

OpenStudy (welshfella):

what is -3-5 ?

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):

Is that the answer? -8?

OpenStudy (tkhunny):

They are so nicely lines up. Think how you would do this. 8 cats + 9 dogs + 14 squirrels Now, subtract 3 squarrels + 3 cats + 2 dogs You wouldn't DREAM of subtracting the dogs from the cats, would you? Dogs just aren't cats. Same things here. x − y = −3 x + 3y = 5 Subtracting, consider the xs separately. consider the ys separately. consider the constants separately Actually, I don't like subtraction. Multiply the bottom one by -1 and add. x − y = −3 -x - 3y = -5 0 - 4y = -8 Much nicer.

OpenStudy (anonymous):

Thank you both!

OpenStudy (tkhunny):

Well, that only finds the value of y. You need the value of x. It would have been smarter to multiply the first equation by 3 and then add Here's where we started. x − y = −3 x + 3y = 5 Multiple the first by 3 3x − 3y = −9 x + 3y = 5 Add 4x + 0 = -4 And it's pretty easy to find x from there.